# UnMarker: A Universal Attack on Defensive Watermarking

### v1

@@ 6.2.2. UnMarker
We now evaluate UnMarker’s ability to universally remove watermarks. In §6.2.3, we compare UnMarker to SOTA attacks and assess its capability of preserving visual similarity.
@@
Configuration. Our analyses and experiments above prove that a conventionally robust scheme (under spatial manipulations) must inject its watermarks in the frequency space since naive manipulations (e.g., cropping) severely impact spatial watermarks that are forced to primarily reside in the background. Schemes that are spatial in nature and therefore rely less on spectral watermarking (e.g., Yu2) may hence evade UnMarker’s spectral optimizations as they are not designed to target these inherently fragile schemes. Yet, equipping UnMarker with a simple cropping layer suffices to defeat such weaker systems. Moreover, as spectral watermarks distribute their modifications throughout the image to minimize visual cues (see §4), cropping also can slightly “weaken” those by eliminating partial modifications residing in the background. This allows using a decreased distortion budget for the optimization stages.
@@
Thus, UnMarker’s design is as follows: We apply mild cropping, followed by high-frequency destructive optimizations (§5.1). Finally, our filtering-based stage (§5.2) eliminates low-frequency traces. This yields a comprehensive attack that targets all carriers, erasing all watermarks. Attackers can adjust the cropping ratio to restrict the output to the actual content. For our experiments, $10\%$ perfectly retains the content. The stages’ order does not affect the results.
@@
While UnMarker relies on several hyperparameters, they should all be calibrated to maximize the spectral loss regardless of the system under attack, complying with our threat model. Instead, they are only affected by properties of the image, such as its size, which, we found to be the only relevant factor. The hyperparameters are in A.3. The size-dependant parameters are merely the thresholds and coefficients. Meanwhile, the number of filters and their kernel sizes are invariant, as by combining multiple filter sizes, a single architecture can defeat any semantic watermark (see 5.2).
@@
Results. We report settings combining different stages of UnMarker, identifying crucial insights that establish our theory. These configurations are CH— the high-frequency destructive process preceded by cropping, CL— the low-frequency optimization (with cropping), HL— the two optimizations without cropping, and CHL, denoting the full UnMarker with all three stages. Results are in Table 2.
@@
Non-Semantic Schemes. Robust non-semantic schemes (w.r.t. known manipulations) cannot alter the content and thus must embed their watermarks in high frequencies. Hence, minimal disruptions to this part of the spectrum where the energy concentration is low suffices to remove them. Our results corroborate this as all such schemes break under CH that combines cropping and high-frequency disruptions. The detection rates of the conventionally robust algorithms (that resist cropping— all but Yu2) remain almost unaffected by CL, which primarily targets low frequencies where they do not operate. Yu2 is defeated by CH, but HL’s results prove this is not due to our optimization since the scheme is insensitive to manipulations in both parts of the spectrum, indicating that cropping only is responsible for this outcome. This comes as no surprise since Yu2 employs spatial watermarking that enables it to resist spectral optimizations but makes it trivially vulnerable to cropping.
Table 3. Comparison with SOTA attacks. For each metric, bold entries correspond to the best attack against the relevant scheme. Gray cells denote scores meeting our quality and complete removal criteria for LPIPS (0.15) and Detect ($50\%$), respectively.
    Yu1 (61%) Yu2 (63%) HiDDeN (73%) PTW (70%) StableSignature (69%) StegaStamp (63%) TRW (0.0141)
@@
UnMarker (CHL)     34.07 0.15 59.08% 31% 22.88 0.1 53.03% 0% 39.56 0.08 58% 0% 11.9 0.14 62.75% 11% 30.87 0.05 52.82% 4% 34.69 0.08 61.49% 43% 60.73 0.1 0.014 40%
  Diffusion Attack no cropping     12.38 0.2 58.31% 28% 13.09 0.17 68.29% 87% 34.1 0.18 59.8% 3% 22.95 0.26 55.53% 7% 36.57 0.32 47.89% 0% 33.15 0.36 83.21% 100% 41.69 0.32 0.0155 99%
+cropping     14.02 0.18 54.24% 8% 11.98 0.17 51.26% 3% 36.97 0.19 56.61% 0% 19.98 0.25 62.01% 16% 41.88 0.31 47.78% 0% 30.83 0.33 77.84% 95% 41.16 0.31 0.0144 79%
  VAE Attack quality=3     36.2 0.15 67.19% 97% 19.67 0.13 73.89% 96% 65.16 0.22 60.4% 8% 26.6 0.26 73.63% 89% 72.21 0.37 48.25% 0% 54.89 0.42 97.9% 100% 71.91 0.38 0.0154 90%
quality=2     46.82 0.18 64.42% 81% 27.4 0.16 66.62% 78% 83.4 0.26 58.36% 1% 39.11 0.29 71.33% 79% 94.85 0.41 44.93% 0% 70.86 0.46 94.7% 100% 93.56 0.41 0.015 88%
quality=1     58.37 0.22 60.63% 59% 38.9 0.19 61% 36% 104.92 0.32 56.26% 2% 54.93 0.33 66.74% 36% 128.88 0.44 43.87% 0% 87.8 0.49 88.84% 100% 118.85 0.46 0.0145 67%
While schemes that process larger images (PTW and StableSignature) are almost insensitive to cropping (see Table 5), they are extremely vulnerable to high-frequency manipulations, evident by their robustness to CL and fragility against HL, bringing their detection to $6\%$ and $9\%$, respectively. Yu1 and HiDDeN that watermark smaller $128{\times}128$ images retain $100\%$ detection even against HL. Yet, their bit accuracies drop sharply from $98.42\%$ and $99.16\%$ (Table 5) to $73.11\%$ and $72.28\%$. The cause for this decline is our high-frequency destruction since CL that combines cropping and low-frequency disruptions offers almost no advantages compared to cropping.
Thus, there is a connection between the image size, sensitivity to cropping, and vulnerability to high-frequency modifications, which is easy to explain: Spatial robustness and stealthiness require spreading the watermark across the image. Meanwhile, non-semantic watermarks are restricted to specific regions (i.e., edges) where they can exploit the high-frequency nature (see §4.1.1) to remain invisible. As the size shrinks, such areas become less prevalent in central parts, forcing the changes to be delegated to the background and making them more vulnerable to cropping. Accordingly, the portion of the change each region receives grows as well, making it difficult to eliminate these watermarks with minimal modifications, explaining why they retain bit accuracies above the threshold despite a significant decline without cropping. Regardless, cropping then becomes capable of eliminating sufficient footprints to further decrease these values below the thresholds, as seen for CH (CHL).
Increasing the size enables schemes to better reject cropping since this new-found resolution creates opportunities for embedding the watermark in central areas. However, the footprint in these regions must be significantly smaller. Thus, minute high-frequency disruptions throughout the image become far more effective, explaining their inability to withstand our high-frequency optimizations alone.
UnMarker (CHL) breaks all non-semantic schemes, with the most resistant being Yu1 that retains a detection rate of $31\%$. Yet, this is worse than a random guess, leaving it useless.
Semantic Schemes. Both StegaStamp and TRW are resistant to high-frequency disruptions, with their detection rates fixed at $100\%$ for CH and the bit accuracies for StegaStamp and inverse distances for TRW retaining almost identical values to those observed when only cropping is applied. Yet, both are extremely vulnerable to low-frequency disruptions, evident by their significant performance decline against CL: StegaStamp’s detection drops to $43\%$ while TRW’s becomes $55\%$. This cements our arguments regarding semantic watermarks being primarily in lower frequencies and the soundness of our filtering-based approach. When comparing CL to the final results (CHL), we observe identical results for StegaStamp, while TRW exhibits an additional decline from $55\%$ to $40\%$, indicating that a small portion of its watermarks is at higher frequencies.
StegaStamp’s watermarks are more visible (see §2.1), allowing them to be engraved in the main content, and increasing their robustness to spatial manipulations (cropping). Since TRW produces less visible watermarks despite targeting low frequencies where significant watermarking footprints are required, it behaves similar to non-semantic schemes for small images, forcing a considerable portion of these changes to populate the background. The results for cropping (Table 5), CL, CH and CHL (Table 2) prove this; while cropping has minimal effects on StegaStamp (the bit accuracy only drops to $93.96\%$), TRW is more sensitive as its inverse distance decreases from $0.022$ to $0.155$ when the threshold is only $0.0141$. Although this is in the detection range, combining cropping with UnMarker’s low-frequency disruptions (CL, CHL) erases its watermarks.
While TRW’s inverse distance is drastically lowered by spectral optimizations (HL), bringing it down to $0.016$ from $0.022$, they do not erase its watermarks, unlike StegaStamp’s. This is because low-frequency distortions affect many pixels simultaneously. Hence, under a strict budget, the number of regions in which these rates can be altered is limited compared to high-frequencies. Thus, StegaStamp, whose watermarks are more concentrated, requires fewer areas to be modified, leading to the optimization being successful without cropping. Yet, we verified that low-frequency disruptions with a higher budget or number of steps eventually break TRW without cropping. However, this may introduce unwanted artifacts unless carefully done. Cropping is as effective but less demanding.
UnMarker defeats all studied systems due to the irreconcilable trade-offs of defensive watermarking, as making watermarks robust to spatial manipulations leaves them vulnerable to spectral disruptions and vice versa. Even when failing to bring the detection to $0\%$, it leaves the scheme useless, with a probability below a random guess.
6.2.3. Comparison with SOTA attacks
Here, we compare UnMarker to regeneration attacks— the VAEAttack (Zhao et al., 2023c; An et al., 2024) and DiffusionAttack(Zhao et al., 2023c; Saberi et al., 2024; An et al., 2024). Despite being certifiably guaranteed to remove non-semantic watermarks, they fail against semantic methods. This robustness creates a false sense of security around semantic watermarks (Zhao et al., 2023c; An et al., 2024), which UnMarker disproves. Additionally, the guarantees ensure non-semantic watermarks will be removed if the attacks introduce sufficient noise into the image, followed by a denoising process to restore it without the watermark that has now been diluted. Stronger watermarks require increased perturbations that prevent perfect reconstruction. Unnatural artifacts can be avoided by using advanced regeneration systems, explaining why the more sophisticated DMs perform better, as they can apply larger distortions while still somewhat accurately retrieving the content, as opposed to off-the-shelf VAEs. Yet, the similarity to the original image may still be lost under such large budgets. UnMarker eliminates all these hurdles, significantly decreasing the distortion while being effective even against the stronger semantic schemes.
Since UnMarker uses cropping, we apply the DiffusionAttack with and without cropping for a fair comparison. The VAEAttack’s image quality was deemed inferior, disqualifying many of its results (see below), and thus a similar comparison was excluded. Results are in Table 3, and samples produced by the three attacks are in A.6.
Visual Quality and Similarity: The range for LPIPS is $[0,1]$. Images with LPIPS distance at the $0.1$ level (i.e., well below $0.2$) are visually similar (Hao et al., 2021). Thus, we choose $0.15$ as a cut-off beyond which we say the attack begins to cause visible changes. This does not mean that slightly higher scores will lead to severe quality degradation since the changes may integrate well with the content. Yet, scores far beyond this threshold attest to the poor nature of the outputs. Despite FID’s known ability to measure similarity, it suffers from two major shortcomings: first, it compares the datasets of watermarked images and those that have undergone the attack and cannot judge samples on an image-by-image basis. Moreover, it may exhibit inconsistencies with human vision, making it less ideal in certain instances (Liu et al., 2018). Thus, while FID is important for assessing quality, we prioritize LPIPS as it can better provide fine-grained judgments.
We do not consider the VAEAttack effective. It allows choosing from eight quality options that control the added noise before regeneration. We apply it with the three lowest options, as we found increasing this factor beyond $3$ makes the attack almost always ineffective. However, all three options severely impact the quality, invalidating the attack. This can be seen in Table 3, as both its FID and LPIPS are extremely worse (larger) than the other attacks whenever it degrades the detection rates. For larger images (over $256{\times}256$), even the best factor ($3$) is destructive, resulting in LPIPS scores at least $3.8x$ compared to UnMarker and far above the $0.15$ cut-off. Its FIDs are also inflated. While its FID against TRW ($71.91$) is close UnMarker’s ($60.73$), its LPIPS is extremely worse, indicating low quality. For smaller images, it retains similarity for Yu1, Yu2 and PTW with $quality=3$, but fails to remove their watermarks. HiDDeN’s watermarks are removed with this factor at an acceptable LPIPS ($0.22$). Still, the failure against all other systems limits this attack’s potential. Lowering the factor impacts quality, invalidating any results for larger images (including PTW). For Yu1 and Yu2, as the attack fails to achieve remarkable removal for $quality=2$, the only relevant case is $quality=1$. The two systems retain acceptable LPIPS scores ($0.22$ and $0.19$) but violate the $0.15$ threshold, and their FID scores are far worse than the other attacks, proving the VAEAttack fails to preserve quality even for these systems (see A.6).
The DiffusionAttack is controlled via a parameter $t$. Following Saberi et al. (Saberi et al., 2024), $t>0.2$ causes significant deviations from the watermarked input, invalidating the attack. We found that for smaller images ($128{\times}128$), this occurs at much lower values. For these images, we set $t=0.05$ as we deemed it the turning point. Lowering $t$ beyond the chosen values results in less successful attacks. UnMarker always achieves better LPIPS. The differences become drastic for larger images (over $256{\times}256$)— $0.1$ at most for UnMarker compared to $\geq 0.37$ for the DiffusionAttack. UnMarker’s LPIPS scores are $0.15$ at most, ensuring indistinguishability from the watermarked content. In contrast, the higher scores obtained by the DiffusionAttack, while not necessarily indicating significant deviations, inevitably entail that some nuances between the image pairs will appear, which may lead to the attack samples being identified as fake, as can be seen in A.6. The attacks’ FIDs alternate, outperforming one another against different schemes. Yet, both attain far better scores than the VAEAttack’s, proving its inferiority.
UnMarker attains the lowest LPIPS, with its worst score being $0.15$, ensuring resemblance to the watermarked images. The similarity between UnMarker’s FIDs and the DiffusionAttack’s, which was previously shown to produce high-quality outputs (Zhao et al., 2023c), further establishes UnMarker’s quality.
Watermark Removal: An attack completely removes watermarks if it brings the detection to $50\%$ or below, as the scheme becomes of no use. Both baselines remove non-semantic watermarks, although the VAEAttack’s quality degradation invalidates its results. UnMarker always outperforms this attack except against StableSignature for which UnMarker brings the detection to $4\%$ while the VAEAttack reduces it to $0\%$. Still, this difference is insignificant. The DiffusionAttack and UnMarker obtain similar results on average, but the DiffusionAttack (+cropping) significantly outperforms UnMarker against Yu1, reducing its detection to $8\%$ compared to $31\%$. Yet, as both drive this rate below $50\%$, specific numbers are irrelevant.

### v2

@@ 6.2.2 UnMarker
We now evaluate UnMarker’s watermark removal capabilities. In §6.2.3, we compare it to SOTA attacks and assess its ability to preserve visual similarity.
@@
Configuration. Our analysis above proves that a conventionally robust scheme (under spatial manipulations) must inject its watermarks in the frequency space. Schemes that are spatial in nature and rely less on spectral watermarking (e.g., Yu2) may hence evade UnMarker’s spectral optimizations as they are not designed to target these inherently fragile schemes. Yet, equipping UnMarker with a simple cropping layer suffices to defeat such weaker systems. Moreover, as spectral watermarks distribute their modifications throughout the image to minimize visual cues (see §4), cropping also can slightly “weaken” those by eliminating partial modifications residing in the background. This allows using a decreased distortion budget for the optimization stages.
@@
Thus, UnMarker’s design is as follows: We apply mild cropping, followed by our high-frequency disruptions (§5.1). Finally, our filtering stage (§5.2) eliminates low-frequency traces. This yields a full attack that targets all carriers, erasing all watermarks. Attackers can adjust the cropping ratio or location to restrict the output to the actual content. For our experiments, $10\%$ center cropping sufficiently retains the content. The stages’ order does not affect the results.
@@
UnMarker’s hyperparameters should be calibrated to maximize the spectral loss regardless of the system under attack, complying with our threat model. Hence, they are only affected by properties of the image, such as its size, which we found to be the only relevant factor. The hyperparameters are in Appendix A. The size-dependant parameters are merely the thresholds and coefficients. Meanwhile, the number of filters and their kernels are invariant, as by combining multiple filter sizes, a single architecture can defeat any semantic watermark (see §5.2).
@@
Results. We report settings combining different stages of UnMarker, identifying crucial insights that establish our theory. These configurations are CH— the high-frequency destructive process preceded by cropping, CL— the low-frequency optimization (with cropping), HL— the two optimizations without cropping, and CHL, denoting the full UnMarker with all three stages. Results are in Table II. Following §6.1, we say that an attack defeats a scheme (alternatively, removes its watermark) if it causes it to lose its Security Advantage (i.e., detection below $50\%$).
@@
Non-Semantic Schemes. Robust non-semantic schemes (w.r.t. known manipulations) cannot alter the content and thus must embed their watermarks in high frequencies. Hence, minimal disruptions to this part of the spectrum where the energy concentration is low suffices to remove them. Our results corroborate this as all such schemes break under CH that combines cropping and high-frequency disruptions. The detection rates of the conventionally robust algorithms (that resist cropping— all but Yu2) remain almost unaffected by CL, which primarily targets low frequencies where they do not operate. Yu2 is defeated by CH, but HL’s results prove this is not due to our optimization since the scheme is insensitive to manipulations in both parts of the spectrum, indicating that cropping only is responsible for this outcome. This comes as no surprise since Yu2 employs spatial watermarking that enables it to resist spectral optimizations but makes it vulnerable to cropping.
Schemes that process larger images (PTW and StableSignature) are hardly sensitive to cropping but vulnerable to high-frequency manipulations, shown by their robustness to CL and fragility against HL, bringing their detection to $6\%$ and $9\%$. Yu1 and HiDDeN that watermark smaller $128{\times}128$ images retain $100\%$ detection against HL. Yet, their bit accuracies drop sharply from $98.42\%$ and $99.16\%$ to $73.11\%$ and $72.28\%$. The cause is our high-frequency destruction since CL that combines cropping and low-frequency disruptions offers almost no advantages over cropping.
Thus, there is a connection between the image size, sensitivity to cropping, and vulnerability to high-frequency modifications, which is easy to explain: Spatial robustness and stealthiness require spreading the watermark across the image. Meanwhile, non-semantic watermarks are restricted to specific regions (i.e., edges) where they can exploit the high-frequency nature (see §4.1.1) to remain invisible. As the size shrinks, such areas become less prevalent in central parts, forcing the changes to be delegated to the background and making them more vulnerable to cropping. Accordingly, the portion of the change each region receives grows as well, making it difficult to eliminate these watermarks with minimal modifications, explaining why they retain bit accuracies above the threshold despite a significant decline without cropping. Regardless, mild cropping then becomes capable of eliminating sufficient footprints to further decrease these values below the thresholds, as seen for CH (CHL).
Larger images enable schemes to better reject cropping since this new-found resolution facilitates embedding the watermark in central areas, but the footprint in these regions must be smaller. Thus, minute disruptions throughout the image become more effective, explaining their inability to withstand high-frequency optimizations alone.
UnMarker (CHL) defeats all non-semantic schemes, with the most resistant being Yu1, which retains a detection rate of $31\%$. Yet, Yu1 still loses its Security Advantage (and so do all other schemes).
Semantic Schemes. StegaStamp and TRW resist high-frequency disruptions (CH), with their detection fixed at $100\%$ and the bit accuracies (inverse distances) retaining almost identical values to those observed for cropping alone. Yet, both are vulnerable to low-frequency disruptions, evident by their significant performance decline against CL: StegaStamp’s detection drops to $43\%$ while TRW’s becomes $55\%$. This cements our arguments regarding semantic watermarks being primarily in lower frequencies and the soundness of our filtering approach. Comparing CL to the final results (CHL), we observe identical detection for StegaStamp. TRW exhibits an additional decline to $40\%$, indicating a small portion of its watermark is in high frequencies.
StegaStamp’s watermarks are more visible (see §2.1), better engraving them in the main content, and increasing their robustness to spatial manipulations (cropping). Since TRW produces less visible watermarks despite targeting low frequencies where significant watermarking footprints are required, it behaves similar to non-semantic schemes for small images, forcing a considerable portion of these changes to populate the background. The results for cropping, CL, CH and CHL prove this; while cropping minimally affects StegaStamp (the bit accuracy only drops to $93.96\%$), TRW is more sensitive as its inverse distance decreases from $0.022$ to $0.155$ when the threshold is only $0.0141$. Although this is in the detection range, combining cropping with UnMarker’s low-frequency disruptions (CL, CHL) erases its watermarks (eliminates its Security Advantage).
Spectral optimizations (HL) drastically lower TRW’s inverse distance (to $0.016$), but do not erase its watermarks. This is because low-frequency distortions affect many pixels simultaneously. Under a strict budget, the number of regions where these rates can be altered is limited compared to high-frequencies. StegaStamp, whose watermarks are centralized, requires modifying fewer areas, making HL successful without cropping. Yet, we verified that low-frequency disruptions with a higher budget or number of steps break TRW without cropping, but this may introduce artifacts unless carefully done. Cropping is as effective but less demanding, confirming a flaw in TRW that targets non-critical regions.
TABLE III: Comparison with SOTA watermark removal attacks.
Note: Performance in the absence of attacks is included for reference. For each metric, bold entries correspond to the best attack against the relevant scheme.
Gray cells denote scores meeting our quality and removal criteria for LPIPS— 0.15 and Detect— $50\%$ (where the system loses its Security Advantage).
(a) Performance against non-semantic schemes.
    Yu1 (61%) Yu2 (63%) HiDDeN (73%) PTW (70%) StableSignature (69%)
@@
None     NA NA 98.42% 100% NA NA 99.71% 100% NA NA 99.16% 100% NA NA 97.96% 100% NA NA 98.97% 100%
  UnMarker (CHL)     34.07 0.15 59.08% 31% 22.88 0.1 53.03% 0% 39.56 0.08 58% 0% 11.9 0.14 62.75% 11% 30.87 0.05 52.82% 4%
  Diffusion Attack no cropping     12.38 0.2 58.31% 28% 13.09 0.17 68.29% 87% 34.1 0.18 59.8% 3% 22.95 0.26 55.53% 7% 36.57 0.32 47.89% 0%
+cropping     14.02 0.18 54.24% 8% 11.98 0.17 51.26% 3% 36.97 0.19 56.61% 0% 19.98 0.25 62.01% 16% 41.88 0.31 47.78% 0%
  VAE Attack quality=3     36.2 0.15 67.19% 97% 19.67 0.13 73.89% 96% 65.16 0.22 60.4% 8% 26.6 0.26 73.63% 89% 72.21 0.37 48.25% 0%
quality=2     46.82 0.18 64.42% 81% 27.4 0.16 66.62% 78% 83.4 0.26 58.36% 1% 39.11 0.29 71.33% 79% 94.85 0.41 44.93% 0%
quality=1     58.37 0.22 60.63% 59% 38.9 0.19 61% 36% 104.92 0.32 56.26% 2% 54.93 0.33 66.74% 36% 128.88 0.44 43.87% 0%
(b) Performance against semantic schemes.
    StegaStamp (63%) TRW (0.0141)
@@
None     NA NA 99.95% 100% NA NA 0.022 100%
  UnMarker (CHL)     34.69 0.08 61.49% 43% 60.73 0.1 0.014 40%
  Diffusion Attack no cropping     33.15 0.36 83.21% 100% 41.69 0.32 0.0155 99%
+cropping     30.83 0.33 77.84% 95% 41.16 0.31 0.0144 79%
  VAE Attack quality=3     54.89 0.42 97.9% 100% 71.91 0.38 0.0154 90%
quality=2     70.86 0.46 94.7% 100% 93.56 0.41 0.015 88%
quality=1     87.8 0.49 88.84% 100% 118.85 0.46 0.0145 67%
UnMarker defeats all semantic schemes due to irreconcilable trade-offs, as robustness to spatial manipulations incurs vulnerability to spectral disruptions and vice versa. Even when not bringing detection to $0\%$, it leaves a scheme useless, eliminating its Security Advantage.
Remarks: Any scheme either modifies the content (semantic) or does not (non-semantic). If attackers know the scheme’s class, they can apply only one of the stages (CH or CL), sufficing to erase the watermark, avoiding additional distortions. This information may be public or obtained by attacking multiple fake images (not the one intended for use in the actual attack), using the two stages independently, and later observing which images were not deleted (flagged). Some schemes may use principles from both classes. For instance, although TRW is semantic, a portion of its watermarks is in high frequencies, as discussed above. Here, applying both stages boosts the success rates. Yet, for all schemes (including less sophisticated ones, where only one stage is required), the full UnMarker (CHL) preserves quality even when both stages are applied (see §6.2.3).
6.2.3 Comparison with SOTA attacks
We compare UnMarker to regeneration attacks— the VAEAttack [35, 31] and DiffusionAttack [35, 29, 31]. Despite being certifiably guaranteed to remove non-semantic watermarks, they fail against semantic methods. This robustness creates a false sense of security around semantic watermarks [35, 31], which UnMarker disproves. Additionally, they guarantee the removal of non-semantic watermarks only if the attack introduces sufficient noise to dilute the watermark, followed by a denoising process to restore the image’s quality. Stronger watermarks require increased perturbations that prevent perfect reconstruction. Unnatural artifacts can be avoided by using advanced regeneration systems, explaining why the more sophisticated DMs outperform VAEs, as they can apply larger distortions while somewhat accurately retrieving the content. Yet, similarity to the original image may still be lost under such large budgets. UnMarker eliminates these hurdles, significantly decreasing the distortion while being effective even against the stronger semantic schemes.
As UnMarker uses cropping, we apply the DiffusionAttack with and without cropping for a fair comparison. The VAEAttack’s image quality was deemed inferior, disqualifying many of its results (see below), and thus a similar comparison was excluded. Results are in Table III, and samples produced by the three attacks are in Appendix E.
Quality: The range for LPIPS is $[0,1]$. Images with LPIPS distance at the $0.1$ level (i.e., below $0.2$) are visually similar [59]. Thus, we choose $0.15$ as a cut-off beyond which we say the attack begins to cause visible changes. This does not mean that slightly higher scores imply severe degradation since the changes may integrate well with the content. Yet, scores far beyond this threshold attest to poor output nature. Despite FID’s known ability to measure quality, it has two major shortcomings: first, it compares the datasets of watermarked images and those that have undergone the attack and cannot judge samples on an image-by-image basis. It may also exhibit inconsistencies with human vision, making it less ideal in certain instances [60]. Thus, while FID is important for assessing quality, we prioritize LPIPS.
We do not consider the VAEAttack effective. It allows choosing from eight quality options that control the added noise before regeneration. We apply it with the three lowest options, as we found increasing this factor beyond $3$ makes the attack almost always ineffective. Yet, all three options severely impact the quality, invalidating the attack. This can be seen in Table III, as both its FID and LPIPS are extremely worse (larger) than the other attacks whenever it degrades the detection rates. For larger images (over $256{\times}256$), even the best factor ($3$) is destructive, resulting in LPIPS scores at least 3.8×3.8$\times$3.8 × compared to UnMarker and well above the $0.15$ cut-off. Its FIDs are also inflated. While its FID against TRW ($71.91$) is close UnMarker’s ($60.73$), its LPIPS is much worse, indicating low quality. For smaller images, it retains similarity for Yu1, Yu2 and PTW with $quality\!\!=\!\!3$, but fails to remove their watermarks. HiDDeN’s watermarks are removed with this factor at an acceptable LPIPS ($0.22$). Still, the failure against other systems limits this attack. Lowering the factor impacts quality, invalidating any results for larger images (including PTW). For Yu1 and Yu2, as the attack fails to achieve remarkable removal for $quality\!\!=\!\!2$, the relevant case is $quality\!\!=\!\!1$. Both systems retain acceptable LPIPS scores ($0.22$ and $0.19$) but violate the $0.15$ threshold, and their FIDs are far worse than the other attacks, proving the VAEAttack fails to preserve quality (see Appendix E).
The DiffusionAttack is controlled via a parameter $t$. Following Saberi et al. [29], $t\!>\!0.2$ causes significant deviations from the watermarked input, invalidating the attack. We found that for smaller images ($128{\times}128$), this occurs at lower values. For these images, we set $t\!\!=\!\!0.05$, deeming it the turning point. A Lower $t$ results in less successful attacks. UnMarker achieves better LPIPS ($\leq\!\!0.15$), ensuring similarity. The differences are drastic for larger images (over $256{\times}256$)— $0.1$ at most for UnMarker compared to $\geq\!0.37$ for the DiffusionAttack. The higher LPIPS of DiffusionAttack, while not necessarily indicating significant deviations, entail that some nuances will appear, which may lead to its outputs being identified as fake— see Appendix E. The attacks’ FIDs alternate, occasionally outperforming one another. Yet, both attain far better FIDs than the VAEAttack.
UnMarker attains the lowest LPIPS, with its worst being $0.15$, ensuring resemblance to the watermarked images. The similarity between UnMarker’s FIDs and the DiffusionAttack’s, which is known to produce high-quality outputs [35], further establishes its capabilities.
Watermark Removal: Both SOTA attacks defeat non-semantic schemes, although the VAEAttack’s quality invalidates its results. UnMarker always outperforms this attack except against StableSignature for which UnMarker brings the detection to $4\%$ while the VAEAttack reduces it to $0\%$. Still, this difference is insignificant. The DiffusionAttack and UnMarker have similar results on average, but the DiffusionAttack (+cropping) significantly outperforms UnMarker on Yu1 ($8\%$ detection vs. $31\%$). Yet, as both eliminate its Security Advantage, specific numbers are irrelevant.
Our results validate regeneration’s known inability to remove semantic watermarks [35, 29, 31]. The VAEAttack moderately affects TRW, especially with quality $1$, bringing its detection to $67\%$, but, as explained above, we consider it invalid. The DiffusionAttack leaves both detection rates almost unchanged despite reducing StegaStamp’s bit accuracy and TRW’s inverse distance from $99.95\%$ and $0.022$ to $83.21\%$ and $0.0155$. While cropping decreases these numbers to $77.84\%$ and $0.0144$ and makes the attack more successful against TRW ($79\%$ detection), UnMarker drastically outperforms it, reducing this to $40\%$ and StegaStamp’s, against which DiffusionAttack is ineffective, to $43\%$.